# A Deep Dive Into How Open-Source Project Maintainers Review and Resolve Bug Bounty Reports

### v1

@@ 5.4.3 Report management
Interviewees argued that having a user-friendly interface and varied features to manage bug reports helped streamline the report-and-resolve process ($I=5$). For instance, one interviewee explained how “being able to link to our security policy, in some ways, the people are looking at that before they’re posting the bug, rather than us having to go back to it afterward, or in some way communicating the expectations for the release cycle.” In this scenario, there is more focus on receiving legitimate reports and clear expectations because of the ability to link a project’s security policy.
@@ 5.5 Security knowledge
Despite the importance of actively practicing secure coding practices and being knowledgeable about general security concerns, most developers do not have a security background. A majority of our study participants, as seen in Table I, also lack a security background. This section reports a series of challenges while conducting bug bounty report reviews and opportunities where OSS project maintainers were able to learn about aspects of security bugs, e.g., CVSS scoring, and how to go about repairing such bugs through hands-on experience.
@@ 5.5.1 Opportunities to learn
Interviewees argued the significance of learning opportunities that happen as a result of bug bounty report review ($I=5$). One interviewee stated, “I think by now I know most attack vectors, so I know a lot of them, and I know I understand with what security researchers come to me with, and I can I have a clear understanding about it.” This reflects the ability of an OSS project maintainer to develop security knowledge by being involved with the bug bounty report review process.
@@ 5.5.2 Knowledge gaps
Interviewees argued the importance of security knowledge gaps hindering bug assessment and resolution ($I=4$). One stated that “we just don’t even have the knowledge nor the expertise nor the time to look into it and all of a sudden we had people, [bug hunters] were reporting security issues, looking into the program, testing it and dedicating their time.”
Some interviewees discuss their experience interacting with security experts and how they were able to “learn a lot [from] just the process alone, even if I didn’t have a security background” ($I=7$). One interviewee believed that being able to shadow someone who reviews bug bounty reports would help them fill in the knowledge necessary to perform effective reviews.
Others discussed the need for starter guides, bug bounty understanding, and security-related FAQs ($I=5$). One interviewee specifically states that they “didn’t get into becoming a maintainer because I was security focused.” In particular, there is concern about setting standards for appropriate behavior from researchers when testing software and interacting with maintainers, understanding the calculation of security metrics such as CVSS scores, and information about CVE creation and qualifications.
@@ 5.5.3 CVE awareness
Some interviewees are under the impression that only bug hunters determine whether a report should receive a CVE ($I=2$). One interviewee described a scenario where a bug hunter “reported what I would consider a security bug with no CVE, I have no idea but that’s just the way they disclosed it, so I was like okay cool.” Another interviewee explained how they did not realize they had a role, or could “have a say”, during CVE assignment. This includes actions such as severity adjustment or updating the description by reaching out to MITRE [6]. Some OSS project maintainers are unaware of such actions.
One interviewee explained, “I’ve had a very miserable experience trying to get MITRE to correct incorrect information on existing CVEs, so like I just kind of avoid dealing with all that.” Although some indicate that they are familiar with the CVE process ($I=6$), others perceive reaching out to CVE authorities as frustrating and actively avoid doing so by making necessary changes to bug reports, i.e., so that there are no mistakes before they are routed for CVE review.
@@ 5.6 Working alongside bug hunters
OSS project maintainers identify that communication and collaboration are crucial to the success of bug bounty report review. Further, although prior work suggests that lack of communication is one of the more challenging components for bug hunters [47, 13, 42, 25], OSS project maintainers felt as if it was one of the least challenging aspects of reviewing bug bounty reports.
@@ 5.6.1 Communication
Several interviewees argued for security expert importance during the review process ($I=7$), two of which have even maintained communication with bug hunters post-disclosure for future assistance with vulnerability finding and repair. We present results on this topic in Section 5.6.2. One interviewee argued that communicating with security experts was important because “they explained to us not just how to reproduce but also the impact,” challenges reported in Section 5.4.1. Another interviewee explained how having access to security experts is needed in OSS bug bounty review because “as maintainers we kind of started at the very beginning, we have so many different hats on and have come into the maintainer-ship through many different avenues,” referencing knowledge gaps reported in Section 5.5.2.
@@
Few interviewees argued that lack of communication is a challenge during bug bounty review ($I=3$). One interviewee described how lack of communication is a problem when they cannot get hunter input for locating the vulnerability in source code or during patch development, both of which details are reported in Section 5.4.2. Another interviewee described unfortunate experiences, initially reported in Section 5.1.2, where bug hunters disclosed vulnerabilities before a patch was released. “I’ve also had times that somebody released one without even talking to me.” While our studies’ results do not emphasize there is a lack of communication, the consequences can be severe and compromise trust in bug bounty programs for OSS.
@@ 5.6.2 Collaboration
Interviewees see bug bounty programs as an avenue to expand their collaboration network ($I=6$). In some cases, the bug hunter becomes an active contributor to the project for which they originally submitted a bug bounty report ($I=2$). “We know that they’re genuinely interested in the improvement of the project, and we really appreciate that kind of contribution.” The same interviewees view bug hunters as contributors, regardless of their decision to continue contributing, but emphasized the importance of their security background and role in helping improve OSS project security posture.
@@ 6.3 Future work
We encourage researchers to investigate further how LLMs can be leveraged to (1) help bug hunters create more effective reports, (2) help reviewers speed up bug bounty review without losing quality and context, and (3) help reviewers generate patches for reported vulnerabilities. We also encourage researchers to develop practical educational resources accessible to developers without a security background that support the bug bounty review process, further investigate other knowledge gaps this subset of developers have about the vulnerability disclosure process, and develop methods to help close such gaps.
@@ 7 Conclusion
In this paper, we conduct a mixed-methods study, i.e., two surveys and semi-structured interviews, to investigate OSS vulnerability disclosure practices by systematically identifying and quantifying the factors that affect how OSS project maintainers’ conduct bug bounty report review. OSS project maintainers find private disclosure and project visibility to be the most beneficial aspects of bug bounty report review, while hunters’ diverted focus and pressure to review bug bounty reports to be the most challenging. We also find that the most helpful features are incentivized reporting and security metric calculation assistance, while collaboration features and additional hunter reputation features are the most wanted capabilities on bug bounty platforms.

### v2

@@ 5.4.3 Report management
Interviewees argued that having a user-friendly interface and varied features to manage bug reports helped streamline the report-and-resolve process ($I=5$). For instance, one interviewee explained how “being able to link to our security policy, in some ways, the people are looking at that before they’re posting the bug, rather than us having to go back to it afterward, or in some way communicating the expectations for the release cycle” (P7). In this scenario, there is more focus on receiving legitimate reports and clear expectations because of the ability to link a project’s security policy.
@@ 5.5 Security knowledge
Despite the importance of actively practicing secure coding practices and being knowledgeable about general security concerns, most developers do not have a security background. A majority of our study participants, as seen in Table I, also lack a security background. This section reports a series of challenges while conducting bug bounty report reviews and opportunities where OSS project maintainers were able to learn about aspects of security bugs, e.g., CVSS scoring, and how to go about repairing such bugs through hands-on experience. Akgul et al. [13] find that bug bounty hunters recognize knowledge gap challenges that could result in report misunderstanding.
@@ 5.5.1 Opportunities to learn
Interviewees argued the significance of learning opportunities that happen as a result of bug bounty report review ($I=5$). One interviewee stated, “I think by now I know most attack vectors, so I know a lot of them, and I know I understand with what security researchers come to me with, and I can I have a clear understanding about it” (P16). This reflects the ability of an OSS project maintainer to develop security knowledge by being involved with the bug bounty report review process.
@@ 5.5.2 Knowledge gaps
Interviewees argued the importance of security knowledge gaps hindering bug assessment and resolution ($I=4$). One stated that “we just don’t even have the knowledge nor the expertise nor the time to look into it and all of a sudden we had people, [bug hunters] were reporting security issues, looking into the program, testing it and dedicating their time” (P6). Some interviewees discuss their experience interacting with security experts and how they were able to “learn a lot [from] just the process alone, even if I didn’t have a security background” ($I=7$). One interviewee believed that being able to shadow someone who reviews bug bounty reports would help them fill in the knowledge necessary to perform effective reviews.
Others discussed the need for starter guides, bug bounty understanding, and security-related FAQs ($I=5$). One interviewee specifically states that they “didn’t get into becoming a maintainer because I was security focused” (P15). In particular, there is concern about setting standards for appropriate behavior from researchers when testing software and interacting with maintainers, understanding the calculation of security metrics such as CVSS scores, and information about CVE creation and qualifications.
@@ 5.5.3 CVE awareness
Some interviewees are under the impression that only bug hunters determine whether a report should receive a CVE ($I=2$). One interviewee described a scenario where a bug hunter “reported what I would consider a security bug with no CVE, I have no idea but that’s just the way they disclosed it, so I was like okay cool” (P13). Another interviewee explained how they did not realize they had a role, or could “have a say” (P17), during CVE assignment. This includes actions such as severity adjustment or updating the description by reaching out to MITRE [6]. Some OSS project maintainers are unaware of such actions.
One interviewee explained, “I’ve had a very miserable experience trying to get MITRE to correct incorrect information on existing CVEs, so like I just kind of avoid dealing with all that” (P10). Although some indicate that they are familiar with the CVE process ($I=6$), others perceive reaching out to CVE authorities as frustrating and actively avoid doing so by making changes to bug reports, i.e., to avoid mistakes before they are routed for CVE review.
@@ 5.6 Working alongside bug hunters
OSS project maintainers identify that communication and collaboration are crucial to the success of report review. Alomar et al. [15] found that bug bounty hunters can build productive relationships with security teams, allowing them to collaborate on fixing vulnerabilities and receiving prompt feedback. Although prior work suggests that lack of communication is one of the more challenging components for bug hunters [52, 13, 46, 28], OSS project maintainers felt as if it was one of the least challenging aspects of reviewing bug bounty reports—confirming opposing views of bug bounty hunters and reviewers.
@@ 5.6.1 Communication
Several interviewees argued for security expert importance during the review process ($I=7$), two of which have even maintained communication with bug hunters post-disclosure for future assistance with vulnerability finding and repair. We present results on this topic in Section 5.6.2. One interviewee argued that communicating with security experts was important because “they explained to us not just how to reproduce but also the impact” (P12), challenges reported in Section 5.4.1. Another interviewee explained how having access to security experts is needed in OSS bug bounty review because “as maintainers we kind of started at the very beginning, we have so many different hats on and have come into the maintainer-ship through many different avenues” (P11), referencing knowledge gaps reported in Section 5.5.2.
@@
Few interviewees argued that lack of communication is a challenge during bug bounty review ($I=3$). One interviewee described how lack of communication is a problem when they cannot get hunter input for locating the vulnerability in source code or during patch development, both of which details are reported in Section 5.4.2. Another interviewee described unfortunate experiences, initially reported in Section 5.1.2, where bug hunters disclosed vulnerabilities before a patch was released. “I’ve also had times that somebody released one without even talking to me” (P9). While our studies’ results do not emphasize there is a lack of communication, the consequences can be severe and compromise trust in bug bounty programs for OSS.
@@ 5.6.2 Collaboration
Interviewees see bug bounty programs as an avenue to expand their collaboration network ($I=6$). In some cases, the bug hunter becomes an active contributor to the project for which they originally submitted a bug bounty report ($I=2$). “We know that they’re genuinely interested in the improvement of the project, and we really appreciate that kind of contribution” (P17). The same interviewees view bug hunters as contributors, regardless of their decision to continue contributing, but emphasized the importance of their security background and role in helping improve OSS project security posture.
@@ 6.3 Future work
We encourage researchers to investigate further how LLMs can be leveraged to (1) help bug hunters create more effective reports, (2) help reviewers speed up reviews without losing quality and context, and (3) help reviewers generate patches for vulnerabilities. One potentially promising direction researchers can explore is harnessing LLMs with content from OSS bug bounty reports to enhance current state-of-the-art LLM techniques used for OSS vulnerability discovery, PoC exploit and test generation, severity determination, and repair. We also encourage researchers to develop practical educational resources accessible to developers without a security background that support the review process, further investigate other knowledge gaps this subset of developers have about the vulnerability disclosure process, and develop methods to help close such gaps.
One multi-dimensional challenge our studies raise is ensuring that fixes proposed by a hunter do not introduce regressions, which is difficult because the CI/CD systems on GitHub are not accessible to or reusable by hunters and may be resolved through automation or infrastructure support so that PoCs can be included as tests into the CI/CD systems of projects. Our Likert study highlights that pressure to review combined with the difficulty of creating test cases for applied fixes, overstated severity, validating a report’s impact, and finding a fix together emphasize the need to automate the severity determination, fixing of vulnerabilities, and verifying the fixes to ease the burden of maintainers. As column 2 of II shows, we found little related work in this space. Automated program repair and efficient regression testing remain open research challenges.
@@ 7 Conclusion
In this paper, we conduct a mixed-methods study, i.e., two surveys and semi-structured interviews, to investigate OSS vulnerability disclosure practices by systematically identifying and quantifying the factors that affect how OSS project maintainers conduct bug bounty report review. OSS project maintainers find private disclosure and project visibility to be the most beneficial aspects of bug bounty report review, while hunters’ diverted focus and pressure to review bug bounty reports to be the most challenging. We also find that the most helpful features are incentivized reporting and security metric calculation assistance, while collaboration features and additional hunter reputation features are the most wanted capabilities on bug bounty platforms.
Appendix A Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
A.1 Summary
This paper investigates the experiences of open-source maintainers on bug bounty platforms through three studies: two surveys and interviews. Based on these studies, the authors identify and categorize 40 characteristics, finding that private disclosure and project visibility are the most important benefits, and provide recommendations for more accommodating bug bounty review processes for open-source maintainers.
A.2 Scientific Contributions
  • Provides a Valuable Step Forward in an Established Field
  • Establishes a New Research Direction
A.3 Reasons for Acceptance
  1. 1.
    This paper provides a valuable step forward in an established field. Namely the perspectives and ranked characteristics of open-source maintainers who conduct bug bounty reviews. A step forward in the research field of software experts’ perspectives concerning security bug bounties.
  2. 2.
    This paper establishes a new research direction. It builds on a series of recent studies investigating motivations for participating in the bug bounty market but considers a new population: open-source maintainers.